OpenStudy (anonymous):

please anyone help me of my assignment. i really nedd further explaination about how getting the answer and i need also help for answering my assignment and make graph every problem these are the probelem: Please anyone can help with my assignment please. I don’t know how to answer this our professor didn’t explain this well PROBLEMS ARE: Give the slope and intercept of each line represented by the reduce equation to the slope- intercept form 1. X-4y=8 2. 7x+y=11 3. 4x-3y=12 4. 3x+7y-6=0 Write the equation of the line determined by the slope m and y- intercept b 1. M=-4, b=7 2. M=-2/3, b=2

OpenStudy (radar):

In 1 thru 4, isolate y using the rules to keep a balanced equation. When you have the y isolated y=mx+b then you have the solution. m, the slope will be the coefficient of x, and b will be the constant or intercept

OpenStudy (radar):

The last 2, are already solved, just write the equation, they have already provided you with the slope m and intercept b!!

OpenStudy (anonymous):

Can you help me to answer some of the problem sir, even the problem of first all problem of first. i need to dwar graph also. i dont know how, im not good at math very much

OpenStudy (radar):

Let us do the first one. The x-4y=8. What about this equation makes it difficult for you?

OpenStudy (radar):

By isolate the y, I mean get the y on the left and the rest of the terms on the right. Like y=???

OpenStudy (radar):

To do this we do algebraic manipulations that does NOT unbalance the equation.

OpenStudy (radar):

Whatever operation is done to the left of the equal an equal operation must be accomplished to the right of the equal sign, if not the equation becomes invalid.

OpenStudy (radar):

Since we were discussing line equation we want our equation to be in the form of: y=mx+b. What we have is x-4y=8. We want to get y only to the left. In this equation y is already on the left so lets get rid of the x on the left. In this case we want to subtract x from the left hand side x-x-4y=????? Since we did something to the left side we must do the same thing on the right, x-x-4y=8-x. Simplify by consolidating similar terms giving us -4y=-x+8. almost there, instead of y alone we have -4y !! What must we do to get just a plain ole positive y? Tune in tomorrow for the exciting transformation of -4y to y!

OpenStudy (radar):

It is tomorrow! Returning to the problem. We were at the point where we had almost isolated y to the left side of our line equation. We have a -4y, we want a y. Whatever we do to the left side, we must do to the right side of the equation. What can we do to the -4y so that we have y. Would adding +5y do it? It doesn't achieve our objective of getting y by itself. Because, we would have to add 5y to the right hand side. Think some more. One way we could do this, is to divide (both sides) by a -4 or multiply both sides by -1/4 either will work. Personally dividing by -4 seems to work better for me: -4y=-x+8 divide both sides by -4 (-4/-4)(y)=(-x/-4)+8/-4 perform the indicated division getting: y=x/4-2 or y=(1/4)(x)-2 This is the desired slope intercept form slope =1/4 intercept =-2 Hope this helped. The rest should be easy for you. Good luck